OpenStudy (anonymous):

what is the value of csc 135?

OpenStudy (loser66):

Still need help?

OpenStudy (anonymous):

yeah :/

OpenStudy (loser66):



OpenStudy (loser66):

135 is a special angle, you should know the values of trig function of it.

OpenStudy (anonymous):

finding sin 135 is easy though, you just use the calculator (.7071067...)? but what is csc?

OpenStudy (loser66):

Reason: \(csc =\dfrac{1}{sin}\)

OpenStudy (loser66):

That's why I need you to find out the value of sin, then take reciprocal of it to get value of csc . dat sit

OpenStudy (anonymous):

so just 1.41421356?

OpenStudy (anonymous):

obviously i will round it lol

OpenStudy (anonymous):

thanks :) that was easier than expected!

OpenStudy (loser66):

\(sin(135) = \dfrac{\sqrt 2}{2}\) , hence, \(csc (135) = \dfrac{2}{\sqrt2}\) or 2

OpenStudy (loser66):

sorry, it is \(\sqrt 2\) , typo

OpenStudy (anonymous):

from my answer choice, it was (square root) 2, but i just square rooted 2 in the calculator and found it to be the same so went with that. thanks for helping me!